Shift report 20030216 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Feb 16 2003 - 16:09:47 EST

  • Next message: Apache: "Shift report 20030216 16:00-24:00"
    ------------------------------------------------------------
    Date: Sun Feb 16 16:09:47 2003, Shift: 08:00-16:00
    Supervisor: AJ, with: SZ
    Log book#14 pages 20 - 21
    
    Shift Summary:
    ------------------------------------------------------------
    1. Machine status during shift
      - 8:15 gap cleaning 
      - 11:30 beams cogged up to now
       
    2. Runs 7964-7966 pedestals
       Many problems with the background ( see e-log)
       one data run: 7981
    3.  T3 triped many times.
        No counting rates in  some runs. 
        Module 1 in T3 off. Flemming and J.H. were to try fix the problems.  
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Sun Feb 16 2003 - 16:13:23 EST